OpenStudy (anonymous):

The decomposition of baking soda (NaHCO3) is given in this reaction: 2NaHCO3 → Na2CO3 + H2O + CO2 If 8 moles of baking soda decompose, how many moles of Na2CO3 are produced? 1 2 4 8

OpenStudy (abhisar):

This can be solved in a similar fashion. First try out yourself. I'll assist you through the process.

OpenStudy (aaronq):

General Scheme for stoichiometric calculations: \(\sf \large 1.\)First write and balance an equation for the process described. \(\sf \large 2.\)Next, use the stoichiometric coefficients to find moles produced. \(\sf \large 3.\) Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\sf \large \color{red}{a}A + \color{blue}{b}B \rightleftharpoons \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\sf \dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\sf\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) \(\sf \large 4.\) Solve